Write the point-slope form of the line satisfying the given conditions. Then use the point-slope form of the equation to write the slope-intercept form of the equation. Passing through (4.12) and (5.15) Type the point-slope form of the equation of the line Simplifty your answer. Use integers or fractions for any number in the equation) Type the slope-intercept form of the equation of the line Simplity your answer Use integers or fractions for any number in the equation.)
Answer #1

( 4, 12 ) and ( 5 , 15 )

slope = ( 15 - 12 ) / ( 5 - 4)

slope = 3

point slope form is

y - 12 = 3 (x - 4 )

slope intercept form is

y = 3x